OpenStudy (anonymous):

limx->-4 (sqrt(x^2+9)-5)/x+4

OpenStudy (anonymous):

the -5 is out of the sqrt

OpenStudy (anonymous):

would the answer be -4-8/-4+4 = 0?

OpenStudy (anonymous):

i mean undefined

OpenStudy (anonymous):

the question is evaluate the limit. no tables allowed

ganeshie8 (ganeshie8):

heard of `rationalizing` before ?

ganeshie8 (ganeshie8):

\[\large \lim\limits_{x\to -4}\dfrac{\sqrt{x^2+9}-5}{x+4}\]

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

oh i see now do i have to multiply by sqrt(x^2+9)+5 top and bottom?

ganeshie8 (ganeshie8):

multiply by 1 : \[\large \dfrac{\sqrt{x^2+9}+5}{\sqrt{x^2+9}+5}\]

ganeshie8 (ganeshie8):

yes you're right!

OpenStudy (anonymous):

ok let me work on it

OpenStudy (anonymous):

so i got 1/sqrt(5)+5 i think i did something wrong

OpenStudy (anonymous):

i think i saw my mistake i forgot to multiply the -5*5

OpenStudy (anonymous):

oh crap i also made a typo i forgot to write x^2

ganeshie8 (ganeshie8):

\[\large \begin{align} \lim\limits_{x\to -4}\dfrac{\sqrt{x^2+9}-5}{x+4}&= \lim\limits_{x\to -4}\dfrac{\sqrt{x^2+9}-5}{x+4} \times \dfrac{\sqrt{x^2+9}+5}{\sqrt{x^2+9}+5} \\~\\&=\lim\limits_{x\to -4} \dfrac{\left(\sqrt{x^2+9}\right)^2-5^2}{\left(x+4\right)\left(\sqrt{x^2+9}+5\right)}\\~\\ &= \lim\limits_{x\to -4} \dfrac{x^2+9-25}{\left(x+4\right)\left(\sqrt{x^2+9}+5\right)}\\~\\ &= \lim\limits_{x\to -4} \dfrac{x^2-16}{\left(x+4\right)\left(\sqrt{x^2+9}+5\right)}\\~\\ &= \lim\limits_{x\to -4} \dfrac{(x-4)(x+4)}{\left(x+4\right)\left(\sqrt{x^2+9}+5\right)}\\~\\ &= \lim\limits_{x\to -4} \dfrac{x-4}{\sqrt{x^2+9}+5}\\~\\ \end{align}\]

ganeshie8 (ganeshie8):

take the limit now

OpenStudy (anonymous):

i got -4/15

OpenStudy (anonymous):

no i meant -4/5